Chapter 8 & 9

k. Know the three phases of Interphase (G1, S, G2), what exactly happens in each phase)

l. What is cytokinesis? After which stage of cell division does it get completed?

m. What is the function of mitosis? (For growth and tissue repair…. explain?)

n. What is the function of meiosis? (For Reproduction….explain?)

o. How many chromosomes do the newly formed cells have at the end of mitosis? is it the same as the parent cell or different?, In other words is it haploid or diploid ?

p. How many chromosomes do the newly formed cells have at the end of meiosis? is it the same as the parent cell or different? In other words is it haploid or diploid?
Answer #1

k.

G1 Phase-Cell grows in size and synthesizes a large amount of proteins especially ribosomes. Mitochondria produce more ATP molecules.

Synthesis (S) Phase-S phase, or synthesis phase, is the period of DNA replication. So the amount of DNA in the cell will be doubled in this phase. Along with DNA synthesis , histones and membrane phospholipids are also synthesized

G2 Phase-the G2 phase cell prepares for mitosis by duplicating the organelles. Synthesizes proteins for mitosis and makes sure that there is no DNA damage or mutation.

L.

Cyotkinesis is cytoplasmic division. It happens after the nuclear division or Karyokinesis. In nuclear division the half of the duplicated chromosomes are packed in each of the newly formed nuclei and this is followed by cytokinesis.

M.

Mitosis divides the cells into two genetically identical copies thereby help in increasing the cell number which in multi cellular organism is required for growth (replacing dead cells and for increasing in size and height) and also in repair of damaged tissue.

N.

Meiosis is required for the production of reproductive cells or gametes. This kind of division occurs in reproductive organs like testes and ovary and produces gamete cells with haploid number (n) of chromosomes.

O.

Newly formed daughter cells will have same number of chromosomes in mitosis. During S phase DNA duplicates (so 2N becomes 4 N even though newly formed chromosomes will be attached to its pair). These attached chromosomes are separated and given to each daughter cells (4N/2 = 2N) during mitosis. So the cells will be diploid.

P.

During meiosis the duplicated genome divides twice. That is 4N/2 = 2N in meiosis I and 2N/2 = N in meiosis II. And four daughter cells are formed each with N number of chromosomes so they all are Haploid.